# A Fundamental Trade-off in Aligned Language Models and its Relation to Sampling Adaptors

### v2

@@ 5.2 Controlling the Trade-off
We can leverage globally normalized sampling adaptors (§ 4) to exercise this control. Globally normalized sampling adaptors directly select strings based on their string probability under the prior. That is, we can show:
${{\widetilde{{p}}}_{+}}({\boldsymbol{y}})$
$\propto{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({{y}_{t}})$
$\propto\ \underbrace{\left(\dots\right){\prod^{|{\boldsymbol{y}}|}_{t=1}}\left(\dots\right)}_{\text{Functions of the Prior, ${p}$}}\,{{p}_{{\textsc{+}}}}({\boldsymbol{y}})$ (13)
where ${{\widetilde{{p}}}_{+}}$ is the induced distribution when applying a transform function ${\boldsymbol{\gamma}}$ to an aligned model ${{p}_{{\textsc{+}}}}$, and ()($\dots$)( … ) refers to the series of scaling and truncation functions that take the prior as input. See App. E for a full derivation.
§ 5.2 says that the effect of applying a globally normalized sampling adaptor to ${{p}_{{\textsc{+}}}}$ is akin to applying it to the prior language model ${p}$ and then multiplying the result by the likelihood that the generated string aligns with human preferences, i.e., the effects of the sampling adaptor can be pushed to the prior. This follows from Eq. 6, Bayes’ rule, and the fact that many transform functions used in globally and locally normalized sampling adaptors can be decomposed into a series of scaling and truncation operations.
Given the probability–quality trade-off, this implies that an appropriate choice of globally normalized sampling adaptor can be used to control the average log-probability of sampled corpora, which then determines the average reward of generated text. For example, we could use the globally normalized equivalent of temperature sampling with a high temperature to produce lower probability (and thus higher reward) strings. We demonstrate this control in § 6.2.
@@ 6.1 A Toy Experiment
The trade-off described in Prop. 1 fundamentally arises as a consequence of typicality and the fact that ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})\propto{p}({\boldsymbol{y}})\exp{{r}}({\boldsymbol{y}})$. We aim to demonstrate these theoretical principles with an easily reproducible toy experiment, where we model these distributions over a finite set of objects ${x}\in{D}$. That is, we will show the existence of the trade-off using the toy models ${{p}_{{\textsc{+}}}}({x})$, ${p}({x})$ and ${{r}}({x})$.
@@
We set $|{D}|=1,000$ and model ${{p}_{{\textsc{+}}}}({x})$ by sampling a distribution from the Dirichlet distribution. Then, we create the prior ${p}({x})$ by applying the softmax to a scaled and noised version of this distribution. That is, we define ${p}(\cdot)\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}\mathrm{softmax}\big{(}{{p}_{{\textsc{+}}}}(\cdot)^{\frac{1}{{\tau}}}+\epsilon)\big{)}$ where $\epsilon\sim U(-\kappa,\kappa)$ and ${\tau}\in{\mathbbm{R}_{>0}}$, $\kappa\in{\mathbbm{R}_{>0}}$ are our hyperparameters. We then define ${{r}}({x})$ analogously to Eq. 4 with ${{r}}({x})=\log\frac{{{p}_{{\textsc{+}}}}({x})}{{p}({x})}$. The distributions of ${{p}_{{\textsc{+}}}}({x})$, ${p}({x})$ and ${{r}}({x})$ over the domain ${D}$ can be seen in App. G. To induce Simpson’s paradox, we tune $\kappa$ and ${\tau}$ such that they are positively correlated, shown in Fig. 1 on the left.
@@
An important point about the trade-off in Prop. 1 is that it occurs with high probability. To illustrate this, we use causal bootstrapping (Little and Badawy, 2020) to construct corpora that are uniformly distributed across 10 bands of average log-probability under the prior. Then, we compute and visualize the corpus probabilities, i.e., $\log{{p}_{{\textsc{+}}}}({\mathcal{X}})$ where ${\mathcal{X}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{\{{x}^{(n)}\}}_{n=1}^{N}$ denotes a corpus of toy objects. If Prop. 1 is correct, we expect to see that corpora exhibiting the trade-off have much higher probability than those that do not. We examine 1,000 corpora sampled this way, each with 100k samples.
@@ 6.2 The Trade-off in Practice
Here we demonstrate the existence of the probability–quality trade-off with an open-sourced aligned language model based on the Llama 2 family (Touvron et al., 2023). Using locally normalized sampling adaptors, we sample a corpus ${\mathcal{Y}}$ of 2,000 texts from an RLHF-tuned model ${q_{{\textsc{+}}}}$. Towards this, we randomly choose 1,000 prompts using the helpfulness dataset from Perez et al. (2022) and for each prompt, we produce two generations. Then, for every string in this corpus, we obtain its log-probability under the prior language model $\log{p}({\boldsymbol{y}})$ and its reward ${{r}}({\boldsymbol{y}})$. The prior and reward models are the same as those used to train ${q_{{\textsc{+}}}}$ in an RLHF scheme. We repeat this using five locally normalized sampling adaptors at five temperatures, totaling 25 sampling schemes and thus $50,000$ $\big{(}\log{p}({\boldsymbol{y}}),{{r}}({\boldsymbol{y}})\big{)}$ pairs. To observe the trade-off, we compute the Pearson and Spearman’s correlation between $\log{p}({\boldsymbol{y}})$ and ${{r}}({\boldsymbol{y}})$ at the string level, and between $\frac{\log{p}({\mathcal{Y}})}{N}$ and $\frac{{{r}}({\mathcal{Y}})}{N}$ at the corpus level.
@@
To compute corpus-level correlations we require a lot of data points of the average log-probability and average reward. However, because sampling multiple corpora is prohibitively expensive, we use standard bootstrap resampling (Efron and Tibshirani, 1994) to create multiple corpora for each of the 25 sampling schemes. Given a corpus of strings generated from ${q_{{\textsc{+}}}}$ with a locally normalized sampling adaptor defined by the transform function ${\boldsymbol{\gamma}}$, we resample uniformly with replacement $N$ times, accepting and rejecting each as described in Algorithm 1. This gives us a resampled corpus ${\mathcal{Y}}^{\prime}$. Then, we compute the average log-likelihood $\frac{\log{p}({\mathcal{Y}}^{\prime})}{N}$ and average reward $\frac{{{r}}({\mathcal{Y}}^{\prime})}{N}$. We do this 2,000 times per sampling scheme, giving us a total of 50,000 $\left(\frac{\log{p}({\mathcal{Y}})}{N},\frac{{{r}}({\mathcal{Y}}^{\prime})}{N}\right)$ pairs, which we then use to we compute the corpus-level correlations. We set $N=200,000$ as preliminary experiments showed that for $N\geq 200,000$ the IMHA converges, i.e, the autocorrelation measured with Cramér’s V falls to $<0.10$.
Sampling adaptors
The five sampling schemes we examine are: top-${k}$ sampling (Fan et al., 2018) for ${k}\in{\{30,50\}}$, nucleus sampling (Holtzman et al., 2020) for ${\pi}\in{\{0.90,0.95\}}$, $\eta$ sampling (Hewitt et al., 2022) and locally-typical sampling (Meister et al., 2023b). For each setting, we examine five temperatures ${\tau}\in{\{0.5,0.75,1.0,1.25,1.5\}}$. This gives us a total of 25 settings that cover various real-world use cases. As a baseline, we also include ancestral sampling.
@@ 7 Results
In both toy and empirical settings, we observe at the corpus level a strong linear anti-correlation between the average log-probability $\frac{\log{p}({\mathcal{Y}})}{N}$ and reward $\frac{{{r}}({\mathcal{Y}})}{N}$. In the toy experiment, as shown in Fig. 1 corpora in the typical set have average log-probabilities and average rewards that exhibit a Pearson correlation of ${r}=-0.92$. And, importantly, these typical corpora occur with significantly higher probability than those that do not. For example, the median log-probability difference between a corpus in and out of the typical set is $10^{84}$ fold.777500,523500523-500,523- 500 , 523 vs. $-500,718$; the difference is somewhat masked by the log-scale. The results in the empirical experiment with real language models are generally consistent with the toy experiment. We observe a Pearson correlation between the average log-probability and average reward of ${r}=-0.93$ with relatively few outliers.
@@
We observe in Fig. 2 that corpora sampled with various globally normalized sampling adaptors are centered at different points on the trade-off and follow qualitatively expected trends. For example, corpora sampled with different temperatures have average log-probabilities that follow the expected order —high temperature corpora have lower average log-probabilities and higher average reward. And, at ${\tau}=1.0$ all sampling adaptors produce corpora with higher average log-probability and lower average reward than ancestral sampling. These results are expected since lower temperature and the examined sampling adaptors skew the sampling distribution towards high probability strings. The behaviour when comparing sampling adaptors with different degrees of truncation also follows expectations, e.g., corpora sampled with nucleus sampling for ${\pi}=0.95$ have lower average reward than those sampled with ${\pi}=0.90$. These findings are in line with our theoretical exposition in § 5.2 and suggest that we can use sampling adaptors to control the average reward of sampled corpora.
@@ 8 Conclusion
There are three main limitations to our work. First, is that we only conduct empirical analysis for English and Transformer-based language models. Second, we don’t experiment over all sampling adaptors, e.g., we did not consider Mirostat-sampling (Basu et al., 2021) or contrastive search decoding (Su et al., 2022) in our experiments. These choices were made because the theory holds independently of these factors, though further work should consider other model architectures, sampling adaptors and models that span a variety of languages and domains. Finally, we have only examined the probability–quality relationship under the paradigm of RLHF (and equivalently, DPO, as we show in App. F), but not other alignment methods like ORPO (Hong et al., 2024) and KTO (Ethayarajh et al., 2024). We leave those to future work.
@@
This paper provides a theoretical analysis of language models. To the best of the authors’ knowledge, this paper has no ethical implications.
@@ Acknowledgments
This research is supported by the Singapore Ministry of Education Academic Research Fund Tier 1 (T1 251RES2216). Josef Valvoda is funded by the Nordic Programme for Interdisciplinary Research Grant 105178 and the Danish National Research Foundation Grant no. DNRF169.
Appendix A Locally-normalized Sampling Adaptors
Despite their empirical success, the normalization performed in locally normalized sampling adaptors can lead to strings that are scored higher by the transform function ${\boldsymbol{\gamma}}$ to have lower probability under the induced language model ${\ddot{{p}}}$. This behaviour stems from the fact that the normalization is dependent on the weights assigned to other tokens at a given time step, and thus leads to inconsistencies at the global level. This is arguably undesirable since it makes it difficult to formally reason about how the transform function—which embodies the core logic of the sampling adaptor—influences the properties of strings sampled from ${\ddot{{p}}}$. To make this clear, we provide an example of this behaviour in top-${k}$ sampling (Fan et al., 2018).
@@ Appendix A Locally-normalized Sampling Adaptors
${p}(aa)=0.5\times 0.4\times 0.4\times 0.4=0.032$ (16)
${p}(bb)=0.5\times 0.4\times 0.4\times 0.3=0.024$ (17)
@@
${\ddot{{p}}}(aa)=0.5\times 0.5\times 0.5\times 0.5=0.0625$ (18)
${\ddot{{p}}}(bb)=0.5\times\frac{0.4}{0.7}\times\frac{0.4}{0.7}\times\frac{0.3}{0.7}=0.06997$ (19)
Because the transform function in top-${k}$ does not modify the token probability if the token is kept (see § E.1), Eq. 16 is precisely the score assigned to the strings by the transform function, and Eq. 18 is the score after the normalization step. And we observe a reversal—the string “$aa$” has a higher score than “$bb$”, but is later assigned a lower probability under the induced model ${\ddot{{p}}}$.
@@ B.2 Proof of the Probability–Quality trade-off
Due to Chebyshev’s inequality, ${\mathbb{P}}\left({\mathcal{Y}}\notin T_{N}^{\varepsilon}({q_{{\textsc{+}}}})\right)\leq\frac{{\mathbb{V}}({I})}{N\varepsilon^{2}}$, we have $\Big{|}C+\frac{\log{p}({\mathcal{Y}})}{N}+\frac{{{r}}({\mathcal{Y}})}{N\beta}\Big{|}<\varepsilon$ with probability at least $\left(1-\frac{{\mathbb{V}}({I})}{N\varepsilon^{2}}\right)$ for all $N$ and $\varepsilon>0$. When $N\geq\frac{{\mathbb{V}}({I})}{\delta\varepsilon^{2}}$, the above holds with probability at least $(1-\delta)$.∎

### v3

@@ 5.2 Controlling the Trade-off
We can leverage sampling adaptors (§ 4) to exercise this control. The intuition for this comes from analyzing the effect of a specific sampling adaptor on the prior string probability. That is, in App. E, we show for particular sampling adaptor that
${{\widetilde{{p}}}_{+}}({\boldsymbol{y}})$ $\propto F_{\boldsymbol{\gamma}}\left[{{p}({\boldsymbol{y}})}\right]{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}({\boldsymbol{y}})\right)$ (13a)
$={\widetilde{{p}}}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}({\boldsymbol{y}})\right),$ (13b)
where ${{\widetilde{{p}}}_{+}}$ is the induced distribution when applying a sampling adaptor to an aligned model ${{p}_{{\textsc{+}}}}$ and $F_{\boldsymbol{\gamma}}\left[{{p}({\boldsymbol{y}})}\right]$ refers to the aggregated series of truncation functions coming from the transform function ${\boldsymbol{\gamma}}$ that rely only on the prior ${p}({\boldsymbol{y}})$. Eq. 13a says that in some special cases, the effect of applying a globally normalized sampling adaptor to ${{p}_{{\textsc{+}}}}$ is akin to applying it to the prior language model ${p}$ and then multiplying the result by the likelihood that the generated string aligns with human preferences, i.e., the effects of the sampling adaptor can be pushed to the prior.
Given the probability–quality trade-off, this suggests that an appropriate choice of sampling adaptor can be used to control the average log-probability of sampled corpora, which then determines the average reward of generated text. For example, we could use the temperature sampling with a high temperature to produce lower probability (and thus higher reward) strings. This intuition proves to be useful and leads to an efficient way to control the trade-off, as we demonstrate in § 6.2.
@@ 6.1 A Toy Experiment
The trade-off described in Prop. 1 fundamentally arises as a consequence of typicality and the fact that ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})\propto{p}({\boldsymbol{y}})\exp(\frac{1}{\beta}{{r}}({\boldsymbol{y}}))$. We aim to demonstrate these theoretical principles with an easily reproducible toy experiment, where we model ${{p}_{{\textsc{+}}}}$ and ${p}$ as language models with support only over a finite subset of ${{{{\Sigma}}^{*}}}$.
@@
Let ${\mathcal{D}}\subset{{{{\Sigma}}^{*}}}$ be a finite set of $|{\mathcal{D}}|=1,000$ strings. [^7] We arbitrarily identify ${\mathcal{D}}$ with ${\{1,2,\ldots 1000\}}$. With this, let us construct a toy aligned language model ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})$ by sampling a distribution over ${\mathcal{D}}$ from the Dirichlet distribution parameterized by a $|{\mathcal{D}}|$-sized vector where every value is set to 0.1. We create the prior ${p}({\boldsymbol{y}})$ by applying the softmax to a scaled and noised version of this distribution. That is, we define ${p}({\boldsymbol{y}})\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}\mathrm{softmax}\big{(}{{p}_{{\textsc{+}}}}(\cdot)^{\frac{1}{{\tau}}}+\epsilon)\big{)}\left({\boldsymbol{y}}\right)$ where $\epsilon\sim U(-\kappa,\kappa)$ and ${\tau}\in{\mathbbm{R}_{>0}}$, $\kappa\in{\mathbbm{R}_{>0}}$ are our hyperparameters. We then define ${{r}}({\boldsymbol{y}})$ analogously to Eq. 4 with ${{r}}({\boldsymbol{y}})=\log\frac{{{p}_{{\textsc{+}}}}({\boldsymbol{y}})}{{p}({\boldsymbol{y}})}$. The distributions of ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})$, ${p}({\boldsymbol{y}})$ and ${{r}}({\boldsymbol{y}})$ over the domain ${\mathcal{D}}$ can be seen in App. G. To induce Simpson’s paradox, we tune $\kappa$ and ${\tau}$ such that they are positively correlated, shown in Fig. 1 on the left.
@@
An important point about the trade-off in Prop. 1 is that it occurs with high probability. To illustrate this, we use causal bootstrapping (Little and Badawy, 2020) to construct corpora that are uniformly distributed across 10 bands of average log-probability under the prior. [^8] This scheme allows us to observe low-probability corpora. Then, we compute and visualize the corpus probabilities, i.e., $\log{{p}_{{\textsc{+}}}}({\mathcal{Y}})$ where ${\mathcal{Y}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{\{{\boldsymbol{y}}^{(n)}\}}_{n=1}^{N}$. Due to Prop. 1, we expect to see that corpora exhibiting the trade-off have much higher probability than those that do not. We examine 1,000 corpora sampled this way, each with 100k samples.
@@ 6.2 The Trade-off in Practice
Here we demonstrate the existence of the probability–quality trade-off with an open-sourced aligned language model based on the Llama 2 family (Touvron et al., 2023). Using locally normalized sampling adaptors, we sample a corpus ${\mathcal{Y}}$ of 2,000 texts from an RLHF-tuned model ${q_{{\textsc{+}}}}$. Towards this, we randomly choose 1,000 prompts using the helpfulness dataset from Perez et al. (2022) and for each prompt, we produce two generations. Then, for every string in this corpus, we obtain its log-probability under the prior language model $\log{p}({\boldsymbol{y}})$ and its reward ${{r}}({\boldsymbol{y}})$. The prior and reward models are the same as those used to train ${q_{{\textsc{+}}}}$ in an RLHF scheme. We repeat this using five locally normalized sampling adaptors at five temperatures, totaling 25 sampling schemes and thus $50,000$ $(\log{p}({\boldsymbol{y}}),{{r}}({\boldsymbol{y}}))$ pairs. To observe the trade-off, we compute the Pearson and Spearman’s correlation between $\log{p}({\boldsymbol{y}})$ and ${{r}}({\boldsymbol{y}})$ at the string level, and between $\frac{\log{p}({\mathcal{Y}})}{N}$ and $\frac{{{r}}({\mathcal{Y}})}{N}$ at the corpus level.
@@
To compute corpus-level correlations we require a lot of data points of the average log-probability and average reward. However, because sampling multiple corpora is prohibitively expensive, we use standard bootstrap resampling (Efron and Tibshirani, 1994) to create multiple corpora for each of the 25 sampling schemes. Given a corpus of strings generated from ${q_{{\textsc{+}}}}$ with a locally normalized sampling adaptor defined by the transform function ${\boldsymbol{\gamma}}$, we resample uniformly with replacement $N$ times, accepting and rejecting each as described in Algorithm 1. This gives us a resampled corpus ${\mathcal{Y}}^{\prime}$. Then, we compute the average log-likelihood $\frac{\log{p}({\mathcal{Y}}^{\prime})}{N}$ and average reward $\frac{{{r}}({\mathcal{Y}}^{\prime})}{N}$. We do this 2,000 times per sampling scheme, giving us a total of 50,000 $\left(\frac{\log{p}({\mathcal{Y}})}{N},\frac{{{r}}({\mathcal{Y}}^{\prime})}{N}\right)$ pairs, which we then use to we compute the corpus-level correlations. We set $N=200,000$ as preliminary experiments showed that for $N\geq 200,000$ the IMHA converges, i.e., the autocorrelation measured with Cramér’s V falls to $<0.10$.
Sampling adaptors.
The five sampling schemes we examine are: top-${k}$ sampling (Fan et al., 2018) for ${k}\in{\{30,50\}}$, nucleus sampling (Holtzman et al., 2020) for ${\pi}\in{\{0.90,0.95\}}$, $\eta$-sampling (Hewitt et al., 2022) and locally-typical sampling (Meister et al., 2023b). For each setting, we examine five temperatures ${\tau}\in{\{0.5,0.75,1.0,1.25,1.5\}}$. This gives us a total of 25 settings that cover various real-world use cases. As a baseline, we also include ancestral sampling.
@@ 7 Results
In both toy and empirical settings, we observe at the corpus level a strong linear anti-correlation between the average log-probability $\frac{\log{p}({\mathcal{Y}})}{N}$ and reward $\frac{{{r}}({\mathcal{Y}})}{N}$. In the toy experiment, as shown in Fig. 1 corpora in the typical set have average log-probabilities and average rewards that exhibit a Pearson correlation of ${r}=-0.92$. And, importantly, these typical corpora occur with significantly higher probability than those that do not. For example, the median log-probability difference between a corpus in and out of the typical set is $10^{84}$ fold.999500,523500523-500,523- 500 , 523 vs. $-500,718$; the difference is somewhat masked by the log-scale. The results in the empirical experiment with real language models are generally consistent with the toy experiment. We observe a Pearson correlation between the average log-probability and average reward of ${r}=-0.93$ with relatively few outliers.
@@
We observe in Fig. 2 that corpora sampled with various globally normalized sampling adaptors are centered at different points on the trade-off and follow qualitatively expected trends. For example, corpora sampled with different temperatures have average log-probability that follow the expected trend—high-temperature corpora have lower average log-probability and higher average reward. And, at ${\tau}=1.0$ all sampling adaptors produce corpora with higher average log-probability and lower average reward than ancestral sampling. These results are expected since lower temperature and the examined sampling adaptors skew the sampling distribution towards high probability strings. The behaviour when comparing sampling adaptors with different degrees of truncation also follows our expectation, e.g., corpora sampled with nucleus sampling for ${\pi}=0.95$ have lower average reward than those sampled with ${\pi}=0.90$. These findings are in line with our theoretical exposition in § 5.2 and suggest that we can use sampling adaptors to control the average reward of sampled corpora.
@@ 8 Conclusion
There are three main limitations to our work. First, is that we only conduct empirical analysis for English and Transformer-based language models. Second, we don’t experiment over all sampling adaptors, e.g., we did not consider Mirostat-sampling (Basu et al., 2021) or contrastive search decoding (Su et al., 2022) in our experiments. These choices were made because the theory holds independently of these factors, though further work should consider other model architectures, sampling adaptors and models that span a variety of languages and domains. Finally, we have only examined the probability–quality relationship under the paradigm of RLHF (and equivalently, DPO, as we show in App. F), but not other alignment methods like ORPO (Hong et al., 2024), ODPO (Amini et al., 2024a) or variational BoN (Amini et al., 2024b). We leave those to future work.
@@
We do not foresee any ethical implications.
@@ Acknowledgments
This research is supported by the Singapore Ministry of Education Academic Research Fund Tier 1 (T1 251RES2216). Josef Valvoda is funded by the Nordic Programme for Interdisciplinary Research Grant 105178 and the Danish National Research Foundation Grant no. DNRF169. Anej Svete is supported by the ETH AI Center Doctoral Fellowship.
Appendix A Locally Normalized Sampling Adaptors
Despite their empirical success, the normalization performed in locally normalized sampling adaptors can lead to strings that are scored higher by the transform function ${\boldsymbol{\gamma}}$ to have lower probability under the induced language model ${\ddot{{p}}}$. This behaviour stems from the fact that the normalization is dependent on the weights assigned to other symbols at a given time step, and thus leads to inconsistencies at the global level. This is arguably undesirable since it makes it difficult to formally reason about how the transform function—which embodies the core logic of the sampling adaptor—influences the properties of strings sampled from ${\ddot{{p}}}$. To make this clear, we provide an example of this behaviour in top-${k}$ sampling (Fan et al., 2018).
@@ Appendix A Locally Normalized Sampling Adaptors
${p}(aaa)=0.5\times 0.4\times 0.4\times 0.4=0.032$ (16)
${p}(bbb)=0.5\times 0.4\times 0.4\times 0.3=0.024$ (17)
@@
${\ddot{{p}}}(aaa)=0.5\times 0.5\times 0.5\times 0.5=0.0625$ (18)
${\ddot{{p}}}(bbb)=0.5\times\frac{0.4}{0.7}\times\frac{0.4}{0.7}\times\frac{0.3}{0.7}=0.06997$ (19)
Because the transform function in top-${k}$ does not modify the symbol probability if the symbol is kept (see § E.1), Eq. 16 is precisely the score assigned to the strings by the transform function, and Eq. 18 is the score after the normalization step. And we observe a reversal—the string “$aa$” has a higher score than “$bb$”, but is later assigned a lower probability under the induced model ${\ddot{{p}}}$.
@@ B.2 Proof of the Probability–Quality trade-off
Due to Chebyshev’s inequality, ${\mathbb{P}}\left({\mathcal{Y}}\notin T_{N}^{\varepsilon}({q_{{\textsc{+}}}})\right)\leq\frac{{\mathbb{V}}({I})}{N\varepsilon^{2}}$, we have $\Big{|}C+\frac{\log{p}({\mathcal{Y}})}{N}+\frac{{{r}}({\mathcal{Y}})}{N\beta}\Big{|}<\varepsilon$ with probability at least $\left(1-\frac{{\mathbb{V}}({I})}{N\varepsilon^{2}}\right)$ for all $N$ and $\varepsilon>0$. When $N\geq\frac{{\mathbb{V}}({I})}{\delta\varepsilon^{2}}$, the above holds with probability at least $1-\delta$.∎